# GO4Align: Group Optimization for Multi-Task Alignment

### v1

@@ Appendix B Algorithm
1:  Input: Maximum iteration number $T$; Batch size $N_{\rm{bz}}$; Learning rate $\alpha$; Temperature hyperparameter $\beta$; Task number $M$; Group number $(1<K\leq M)$; Calibration vector $\bm{\mathcal{Q}}(\bm{\theta})$.
@@
3:  Initialize each calibration value $q^{m}_{0}\leftarrow{\frac{1}{M}}$;
4:  for $t=0:T$ do
@@
5:     Randomly sample a batch of training samples: $\{({\bm{x}}_{n},y_{n}^{1},...,y_{n}^{M})\}_{n=1}^{N_{\rm{bz}}},\text{where}~{}\bm{x}_{n}\in\mathcal{X},y_{n}^{m}\in\mathcal{Y}^{m};$
@@
6:     Compute the empirical risk for each task at the $t$-th iteration: ^$\hat{\mathcal{L}}^{m}(\bm{\theta}^{s}_{t},\bm{\theta}^{m}_{t})=\frac{1}{N_{\rm{bz}}}\sum_{n=1}^{N_{\rm{bz}}}\ell^{m}(f({\bm{x}}_{n};\bm{\theta}^{s}_{t},\bm{\theta}^{m}_{t}),y_{n}^{m});$ // Lower-level optimization for grouping-based task interactions.
7:     Compute the scale vector ${{\mathcal{P}}_{t}(\bm{\theta}_{t})}$ in Eq. (5);
8:     Compute the smoothness vector ${{\mathcal{Q}}_{t}(\bm{\theta}_{1:t})}$ in Eq. (6);
9:     Obtain the group indicators ${\gamma}_{t}(\bm{\theta}_{1:t})$ with Eq.(7);
@@
10:     Update the group information $\bm{\omega}_{t}$ and $\bm{\mathcal{G}}_{t}$ with Eq. (3) and the obtained ${\gamma}_{t}(\bm{\theta}_{1:t})$. // Upper-level optimization for model parameters.
@@
11:     Update the model parameters $\bm{\theta}$ with Eq. (4): $\bm{\theta}_{t+1}\leftarrow\operatorname*{argmin}_{\bm{\theta}}{{{\bm{\omega}}_{t}}^{\top}}{\bm{\mathcal{G}}}_{t}\hat{\bm{L}}(\bm{\theta}).$
12:  end for
Appendix C Experimental Set-Up & Implementation Details
C.1 Benchmark Descriptions
@@
NYUv2 Nathan Silberman & Fergus (2012) is an indoor scene dataset consisting of 1449 RGBD images and dense per-pixel labeling with 13 classes. The learning objectives include $3$ different dense prediction tasks: image segmentation, depth prediction, and surface normal prediction based on any scene image.
@@
CityScapes Cordts et al. (2016) contains 5000 street-view RGBD images with per-pixel annotations. It needs to predict $2$ dense prediction tasks: image segmentation and depth prediction.
@@
QM9 Blum & Reymond (2009) is a benchmark for group neural networks to predict $11$ properties of molecules. It consists of $>$130K molecules represented as graphs annotated with node and edge features. We use $110K$ molecules from the QM9 example in PyTorch Geometric Fey & Lenssen (2019), $10K$ molecules for validation, and the rest of $10K$ molecules as a test set.
@@
CelebA Liu et al. (2015) contains 200K face images of 10K different celebrities, and each face image is provided with 40 facial binary attributes. As the protocol provided in previous work Liu et al. (2023), each attribute corresponds to one task. Thus, we consider CelebA as a 40-task MTL problem.
C.2 Compared Multi-task Learning Baselines
@@
From the gradient manipulation branch, (1) MGDA Sener & Koltun (2018) that finds the equal descent direction for each task; (2) PCGrad Yu et al. (2020) proposes to project each task gradient to the normal plan of that of other tasks and combining them together in the end; (3) CAGrad Liu et al. (2021) optimizes the average loss while explicitly controls the minimum decrease across tasks; (4) IMTL-G Liu et al. (2020) finds the update direction with equal projections on task gradients; (5) GradDrop Chen et al. (2020) that randomly dropout certain dimensions of the task gradients based on how much they conflict; (6) NashMTL Navon et al. (2022) formulates MTL as a bargaining game and finds the solution to the game that benefits all tasks.
@@
From the loss scaling branch, (1) Linear scalarization (LS) is the sum of empirical risk minimization; (2) Scale-invariant (SI) is invariant to any scalar multiplication of task losses; (3) Dynamic Weight Average (DWALiu et al. (2019), a heuristic for adjusting task weights based on rates of loss changes; (4) Uncertainty Weighting (UWKendall et al. (2018) uses task uncertainty as a proxy to adjust task weights; (5) Random Loss Weighting (RLWLin et al. (2021a) that samples task weighting whose log-probabilities follow the normal distribution; (6) FAMO Liu et al. (2023) decreases task losses approximately at equal rates.
C.3 Neural Architectures & Training Details
@@
For NYUv2 and CityScapes, we follow the training and evaluation protocol in Navon et al. (2022), which adds data augmentations during training for all compared methods. We train each method for 200 epochs with an initial learning rate of $41e-$ 4 and reduce the learning rate to $55e-$ 5 after 100 epochs. The architecture is Multi-Task Attention Network (MTAN) Liu et al. (2019) built upon SegNet Badrinarayanan et al. (2017). Batch sizes for NYUv2 and CityScapes are set as 2 and 8 respectively. To make a fair comparison with previous works Liu et al. (2019, 2021, 2023); Yu et al. (2020), we report the test performance averaged over the last 10 epochs.
@@
We follow the protocol in  Navon et al. (2022) to normalize each task target at the same scale for fairness. We train each method for 300 epochs with a batch size of 120 and search for the best learning rate in ${\{1e-3,5e-4,1e-4\}}$. We take ReduceOnPlateau Navon et al. (2022) as the learning-rate scheduler to decrease the lr once the validation overall performance stops improving. The validation set is also used for early stopping.
@@
Following Liu et al. (2023), we use a neural network with five convolutional and two fully connected layers as the shared encoder. The decoder of each task is implemented by another fully connected layer. We train the model for 15 epochs with a batch size of 256. We adopt Adam as the optimizer with a fixed learning rate of $31e-$ 3. Similar to QM9, we use the validation set for early stopping and hyperparameter selection, such as the number of groups $K$ and the step size of smoothness value $\beta$. We conduct all experiments on a single NVIDIA A100 GPU.
Appendix D Additional Experimental Results
D.1 Detailed results on QM9 and CityScapes
@@
D.2 Training time comparisons
In Fig. 4, we show the training time of MTO methods relative to a baseline method (LS). Here we provide real training time (seconds) in Table 6, where we compute the average training time of randomly 5 epochs. From this table, we observe that loss-oriented methods in general use less time for one epoch than gradient-oriented methods. GO4Align requires the second lowest time cost during training on all four datasets, demonstrating its good computational efficiency.
@@ D.2 Training time comparisons
D.3 Visualizations on risk ratios
@@
D.4 Dynamical group assignment with different group indicators
To tractably solve the bi-level problem in the introduced adaptive group risk minimization principle, we decompose the whole optimization process into two entangled phases: (i) dynamical group assignment and (ii) risk-guided group indicator. In addition to the designed group indicator, the adaptive group risk minimization can also be plug-and-play in previous methods by utilizing their learnable task weights as the group indicators.
The results of previous methods (MGDA Sener & Koltun (2018), NashMTL Navon et al. (2022), and FAMO Liu et al. (2023)) with the proposed adaptive group risk minimization are reported in Table 7. The experiments are conducted on NYU-v2. As shown in the table, when combined with the adaptive group risk minimization, the previous methods further improve the overall performance. MGDA with adaptive group risk minimization achieves the most obvious improvement. Moreover, our method still outperforms the other methods. The reason can be our group indicator considers both the inter-task and historical information, whereas the task scales of other methods do not.
MGDA 13.25 -9.11 0.83 1.38
MGDA+Sec.3.2(Ours) 6.06 -11.69 -1.14 -1.89
NashMTL -4.09 -22.01 3.15 -4.04
NashMTL+Sec.3.2(Ours) -7.75 -20.14 3.60 -4.20
FAMO -1.65 -20.02 1.29 -4.10
FAMO +Sec.3.2(Ours) 1.76 -21.17 -0.03 -4.32
Sec.3.3(Ours)+Sec.3.2(Ours) -4.03 -20.37 -1.18 -6.08
Table 7: Adaptive group risk minimization with different group indicators. With the adaptive group risk minimization, previous methods perform better on the overall results. Moreover, our method still performs the best compared with the others.

### v2

@@ Appendix A Algorithm of GO4Align
1:  Input: Maximum iteration number $T$; Batch size $N_{\rm{bz}}$; Learning rate $\alpha$; Temperature hyperparameter $\beta$; Task number $M$; Group number $(1<K\leq M)$;
@@
3:  for $t=0:T$ do
@@
4:     Randomly sample a batch of training samples: $\{({\bm{x}}_{n},y_{n}^{1},...,y_{n}^{M})\}_{n=1}^{N_{\rm{bz}}},\text{where}~{}\bm{x}_{n}\in\mathcal{X},y_{n}^{m}\in\mathcal{Y}^{m};$
@@
5:     Compute the empirical risk for each task at the $t$-th iteration: ^$\hat{\mathcal{L}}^{m}(\bm{\theta}^{s}_{t},\bm{\theta}^{m}_{t})=\frac{1}{N_{\rm{bz}}}\sum_{n=1}^{N_{\rm{bz}}}\ell^{m}(f({\bm{x}}_{n};\bm{\theta}^{s}_{t},\bm{\theta}^{m}_{t}),y_{n}^{m});$ // Lower-level optimization for grouping-based task interactions.
6:     Compute the scale vector ${{\mathcal{P}}_{t}(\bm{\theta}_{t})}$ in Eq. (5);
7:     Compute the smoothness vector ${{\mathcal{Q}}_{t}(\bm{\theta}_{1:t})}$ in Eq. (6);
8:     Obtain the group indicators ${\bm{\gamma}}_{t}(\bm{\theta}_{1:t})$ with Eq.(7);
@@
9:     Update the group information $\bm{\omega}_{t}$ and $\bm{\mathcal{G}}_{t}$ with Eq. (3) and the obtained ${\bm{\gamma}}_{t}(\bm{\theta}_{1:t})$. // Upper-level optimization for model parameters.
@@
10:     Update the model parameters $\bm{\theta}$ with Eq. (4): $\bm{\theta}_{t+1}\leftarrow\operatorname*{argmin}_{\bm{\theta}}{{{\bm{\omega}}_{t}}^{\top}}{\bm{\mathcal{G}}}_{t}\hat{\bm{L}}(\bm{\theta}).$
11:  end for
Figure 7: Optimization process of the proposed adaptive group risk minimization principle. At each iteration, given the randomly sampled mini-batch data, we first compute the group information $\bm{\omega}$ and $\bm{\mathcal{G}}$ in the lower-level optimization and then update the model’s parameter $\bm{\theta}$ in the upper-level optimization.
Appendix B Experimental Set-Up & Implementation Details
B.1 Benchmark Descriptions
@@
NYUv2 [34] is an indoor scene dataset consisting of 1449 RGBD images and dense per-pixel labeling with 13 classes. The learning objectives include $3$ different dense prediction tasks: image segmentation, depth prediction, and surface normal prediction based on any scene image.
@@
CityScapes [35] contains 5000 street-view RGBD images with per-pixel annotations. It needs to predict $2$ dense prediction tasks: image segmentation and depth prediction.
@@
QM9 [36] is a benchmark for group neural networks to predict $11$ properties of molecules. It consists of $>$130K molecules represented as graphs annotated with node and edge features. We use $110K$ molecules from the QM9 example in PyTorch Geometric [44], $10K$ molecules for validation, and the rest of $10K$ molecules as a test set.
@@
CelebA [37] contains 200K face images of 10K different celebrities, and each face image is provided with 40 facial binary attributes. As the protocol provided in previous work [14], each attribute corresponds to one task. Thus, we consider CelebA as a 40-task MTL problem.
This work follows the same experimental setting used in NashMTL [13] and FAMO [14], including the dataset partition for training, validation, and testing. The benchmark partition is shown in Table 6. We also note that NYUv2 and Cityscapes do not have validation sets. Following the protocol in  [13, 14], we report the test performance averaged over the last ten epochs.
Table 6: Benchmark partition for training, validation, and testing.
Datasets Total Training Validation Test
NYUv2 1449 795 N/A 654
CityScapes 3475 2975 N/A 500
QM9 $\sim$130k $\sim$110k 10k 10k
CelebA 202,599 162,770 19,867 19,962
B.2 Compared Multi-task Learning Baselines
@@
From the gradient manipulation branch, (1) MGDA [18] that finds the equal descent direction for each task; (2) PCGrad [12] proposes to project each task gradient to the normal plan of that of other tasks and combining them together in the end; (3) CAGrad [16] optimizes the average loss while explicitly controls the minimum decrease across tasks; (4) IMTL-G [38] finds the update direction with equal projections on task gradients; (5) GradDrop [17] that randomly dropout certain dimensions of the task gradients based on how much they conflict; (6) NashMTL [13] formulates MTL as a bargaining game and finds the solution to the game that benefits all tasks.
@@
From the loss scaling branch, (1) Linear scalarization (LS) is the sum of empirical risk minimization; (2) Scale-invariant (SI) is invariant to any scalar multiplication of task losses; (3) Dynamic Weight Average (DWA[6], a heuristic for adjusting task weights based on rates of loss changes; (4) Uncertainty Weighting (UW[5] uses task uncertainty as a proxy to adjust task weights; (5) Random Loss Weighting (RLW[39] that samples task weighting whose log-probabilities follow the normal distribution; (6) FAMO [14] decreases task losses approximately at equal rates.
B.3 Neural Architectures & Training Details
@@
For NYUv2 and CityScapes, we follow the training and evaluation protocol in [13], which adds data augmentations during training for all compared methods. We train each method for 200 epochs with an initial learning rate of $41e-$ 4 and reduce the learning rate to $55e-$ 5 after 100 epochs. The architecture is Multi-Task Attention Network (MTAN) [6] built upon SegNet [45]. Batch sizes for NYUv2 and CityScapes are set as 2 and 8 respectively. To make a fair comparison with previous works [6, 16, 14, 12], we report the test performance averaged over the last 10 epochs.
@@
We follow the protocol in  Navon et al. [13] to normalize each task target at the same scale for fairness. We train each method for 300 epochs with a batch size of 120 and search for the best learning rate in ${\{1e-3,5e-4,1e-4\}}$. We take ReduceOnPlateau [13] as the learning-rate scheduler to decrease the lr once the validation overall performance stops improving. The validation set is also used for early stopping.
@@
Following [14], we use a neural network with five convolutional and two fully connected layers as the shared encoder. The decoder of each task is implemented by another fully connected layer. We train the model for 15 epochs with a batch size of 256. We adopt Adam as the optimizer with a fixed learning rate of $31e-$ 3. Similar to QM9, we use the validation set for early stopping and hyperparameter selection, such as the number of groups $K$ and the step size of smoothness value $\beta$. We conduct all experiments on a single NVIDIA A100 GPU.
Appendix C Additional Experimental Results
C.1 Detailed Results on QM9 and CityScapes
@@
C.2 Training Time Comparisons
In Fig. 4, we show the training time of MTO methods relative to a baseline method (LS). Here we provide real training time (seconds) in Table 9, where we compute the average training time of 5 epochs. From this table, we observe that loss-oriented methods in general use less time for one epoch than gradient-oriented methods. GO4Align requires the second lowest time cost during training on all four datasets, demonstrating its good computational efficiency.
@@ C.2 Training Time Comparisons
C.3 Visualizations on Risk Ratios
@@
Appendix D Additional Related Work
In MTL, we group multi-task learning methods into three categories: multi-task optimization, multi-task grouping, and deep multi-task architecture. The first two categories are most related to this paper and are mentioned in the main paper. To be self-contained, here we provide detailed discussions about deep multi-task architecture.
Our method GO4Align builds on multi-task grouping and multi-task optimization, inheriting advantages from both to balance different tasks in joint learning. In the following, we discuss each category of multi-task learning methods and explain how they relate to our proposed method.
Deep Multi-Task Architecture. Multi-task architecture design can be roughly categorized into either a hard-parameter sharing design [46, 47] or a soft-parameter sharing design [48, 49, 6]. The hard-parameter sharing design generally contains a shared encoder and several task-specific decoders. Branching points between the shared encoder and decoders are determined in an ad-hoc way [1, 50], resulting in a suboptimal solution. Some work [51, 47, 52] automatically learns where to share or branch with a network. The soft-parameter sharing design considers all parameters task-specific and instead learns feature-sharing mechanisms to handle the cross-task interactions [48, 53]. Some work [54, 55] are probabilistic multi-task learning methods, which introduce latent variables to encode task-specific information and enable knowledge sharing in the latent spaces. Soft-parameter sharing methods usually struggle with model scalability as the model size grows linearly with the number of tasks. GO4Align focuses on the optimization of multi-task learning, which adopts a simple hard-parameter sharing architecture as the backbone and subsequently reduces task interference in the gradient space of the shared encoder or the loss space.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The contributions and scope of this paper are claimed in the abstract. Detailed information can be found in the introduction section 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We provide a "limitation" subsection in the conclusion section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide all the details of the experiment in the experiment section. The algorithm and Python codes of our method can be found in Appendix A.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide all experimental details in the experiment section. The algorithm and Python codes of our method can be found in Appendix A. Details about MTL benchmarks in this paper are provided in Appendix B.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We follow the standard experimental setup in MTO, where the data splits, hyperparameters, and optimizer are set as the same as previous works (nashMTL, CAGrad, and FAMO). Detailed information can be found in Appendix B.3.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Following the standard experimental setup, we repeat each experiment over 3 random seeds and report the mean of the results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the computing resources in Appendix B.3. The training-time efficiency is provided in Figure 4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We reviewed and followed the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We provide the potential broader impacts in the conclusion section 6.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The data and models pose no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the original papers that produced the code package and datasets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The provided Python code cannot be used without the authors’ permission.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.